OpenStudy (anonymous):

A fireworks company has two types of rockets called Zinger 1 and Zinger 2. The polynomial -16t2 + 150t gives the height in feet of Zinger 1 at t seconds after launch. The polynomial -16t2 + 165t gives the height of Zinger 2 at t seconds after launch. If the rockets are launched at the same time and both explode 6 seconds after launch, how much higher is Zinger 2 than Zinger 1 when they explode?

OpenStudy (hunus):

What is the height of both of the rockets at t = 6?

OpenStudy (anonymous):

sorry tring to wrap my head around this one.

OpenStudy (anonymous):

t= time?

OpenStudy (hunus):

Yup

OpenStudy (hunus):

Here's a picture of their graphs

OpenStudy (anonymous):

zinger 1= 406, zinger 2 = 421

OpenStudy (anonymous):

Id have to say 90 secounds beased on the answers given to me.

OpenStudy (anonymous):

I mean feet

OpenStudy (hunus):

You're around the right numbers \[Z_{1}=-16t^2+150t\] \[Z_{2}=-16t^2+165t\] so at t=6 we will have \[Z_{1}=-16*6^2+150*6 = -16*36+900 = 900 - 576= 324\]

OpenStudy (hunus):

That's a good guess. Try calculating zinger 2's height and seeing their heigh difference

OpenStudy (anonymous):

lol i was way off again. let me re read that see if I can't figure that out.

OpenStudy (anonymous):

where did you get 900 from?

OpenStudy (hunus):

150*6

OpenStudy (anonymous):

oh okay gotchya.

OpenStudy (hunus):

Did you get it?

OpenStudy (anonymous):

I did thankyou